# Prompting Underestimates LLM Capability for Time Series Classification

### v1

@@ 1 Introduction
LLMs demonstrate broad generalization across natural language tasks, supporting flexible pattern matching, contextual reasoning, and symbolic manipulation. These capabilities have motivated efforts to extend LLMs to non-linguistic data Fons et al. (2024); Mai et al. (2024). Recent work has explored LLMs for time-series forecasting, classification, and analysis Jain et al. (2023); Gruver et al. (2024); Merrill et al. (2024b); Morales-García et al. (2024), often relying on major modifications and retraining. However, strong end-to-end performance in these settings does not necessarily imply that LLMs encode discriminative temporal structure, as results may instead reflect task formulation, additional fine-tuning, or prompting choices.
@@
Merrill et al. (2024a) show that, despite some success in zero-shot forecasting, LLMs struggle to reliably extract discriminative information from time series, performing near-randomly on reasoning and question-answering tasks that require interpreting temporal structure. Others similarly show that removing the LLM component from time-series pipelines often does not degrade performance and can even improve it, calling into question the necessity of language models for these tasks Tan et al. (2024). In contrast, a growing body of work argues that LLM-centric approaches may enable more general and interactive forms of time series understanding that extend beyond task-specific prediction Jin et al. (2024b). Together, these mixed findings raise a central question: do observed limitations reflect a lack of time-series understanding in LLMs, or a mismatch between what these models represent and how that information is extracted?
@@ 2 Related Work
Probing and Hidden Representations in LLMs. Early work on Layer-wise Relevance Propagation (LRP) introduced methods for attributing model predictions to intermediate representations Bach et al. (2015) which later extended to more complex architectures Binder et al. (2016); Arras et al. (2017). Building on this work, probing methods emerged to explicitly measure what information these representations encode by testing their linear separability with lightweight classifiers Alain and Bengio (2018).
In recent years, probing methods have increasingly been used as to extract features for a wide range of downstream applications. Zhu et al. (2025) and Wang et al. (2025), show that hidden states can be leveraged to estimate the difficulty an LLM is anticipated to have on a query prior to generation. More broadly, prior work has demonstrated that hidden representations encode rich and actionable signals that can be exploited across diverse settings, including alignment Kong et al. (2024); Li et al. (2023); Zhang et al. (2025), safety Wang et al. (2024); Lu et al. (2025), interpretability Ghandeharioun et al. (2024); Jacobi and Niv (2025), and robustness Lad et al. (2025); Yan et al. (2024).
@@
Time-Series Research. Time series research spans a wide range of objectives, with forecasting and imputation forming the historical foundation of the field. Classical statistical methods, such as ARIMA and Kalman filters, model linear temporal dependencies Box et al. (2015); Hamilton (1994); Hyndman and Athanasopoulos (2018), while neural architectures, including LSTMs, convolutional networks, and transformers, enable richer nonlinear and long-range temporal modeling Hochreiter and Schmidhuber (1997); Bai et al. (2018); Zhou et al. (2021); Wu et al. (2021); Lim et al. (2021).
Building on this progression, recent work explores how large language models can be adapted for time series analysis. Existing approaches largely follow a small number of recurring patterns Zhang et al. (2024): direct prompting that treats time series as text, often with modified numerical tokenization Gruver et al. (2024); Xue and Salim (2024); Liu et al. (2023); Xie et al. (2023); discretization or quantization schemes that map signals to token sequences Chung et al. (2023); Duan et al. (2023); Yang et al. (2024); Rubenstein et al. (2023); Yu et al. (2023); and architectures that introduce dedicated time-series encoders aligned with language model representations King et al. (2023); Sun et al. (2024); Zhou et al. (2023c); Qiu et al. (2023); Zhou et al. (2023a); Chang et al. (2025); Jin et al. (2024a). Related work also explores visual time-series representations processed by vision–language models or uses LLMs as controllers for external analysis tools Girdhar et al. (2023); Su et al. (2023); Moon et al. (2023); Zhang et al. (2023).
@@ 5 Conclusion
Acknowledgements
This material is based upon work supported by the National Science Foundation (NSF) under Grant No. 2145357. This research was also partially sponsored by the Army Research Laboratory and was accomplished under the Cooperative Agreement Number W911NF-24-2-0180. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Laboratory or the U.S. Government.
@@ A.1 Baselines
Attend follows the dedicated time-series encoder family. It uses self-attention to model cross-channel and temporal dependencies in multivariate sensor data and learns discriminative representations directly from raw time series, without adapting language-model token spaces Abedin et al. (2021). InstructTime belongs to the time-series quantization and instruction-based prompting family. It discretizes time series into vector-quantized tokens, aligns them with domain-specific textual instructions, and uses generative instruction tuning to produce label text via next-token prediction Cheng et al. (2025). OneFitsAll follows the encoder-alignment family. It reuses a frozen pretrained transformer from language or vision domains and learns lightweight input projections and normalization layers to map time-series patches into the model’s embedding space, enabling a single backbone across tasks Zhou et al. (2023b). TS2Vec represents dedicated time-series representation learning. It learns multi-scale contextual embeddings through hierarchical contrastive learning over augmented temporal views, which are aggregated and used by simple downstream classifiers without language models Yue et al. (2022). Moment belongs to the TS-Encoder family. It is a high-capacity transformer encoder pre-trained on large-scale, multi-domain time series using masked time-series modeling, and serves as a general-purpose representation learner that can be adapted to classification, forecasting, anomaly detection, and imputation via lightweight task-specific heads or linear probing (Goswami et al., 2024).
@@ A.2 Models
To assess the generality of our findings, we evaluate a diverse set of pretrained models spanning different architectures and scales. [^1] All models are open source, as access to internal hidden states is required for probing. Motivated by prior work showing that visualizations provide a strong representation for time series classification Liu et al. (2025), we primarily focus on vLLMs. Specifically, we evaluate Llama-3.2-11B-Vision-Instruct 71, Qwen/Qwen2.5-VL-32B-Instruct Bai et al. (2025), and Mistral-Small-3.1-24B-Instruct-2503 AI (2025). To verify that our conclusions extend beyond multimodal architectures, we also evaluate two text-only LLMs: GPT-oss OpenAI et al. (2025) and google/gemma-2-9b-it Team et al. (2024).
@@ A.3 Datasets
CTU records 24-hour household electricity usage patterns and classifies computers as desktops or laptops, capturing differences in device power-consumption cycles Chen et al. (2015). EMG consists of muscle response recordings under electrical stimulation, categorized into healthy, myopathy, and neuropathy classes Goldberger and others (2000). HAR contains multivariate accelerometer recordings of human activities such as walking, sitting, and stair climbing, and exhibits substantial inter-subject variability Anguita et al. (2013). HAD similarly contains multivariate human activity data; we use both the tri-axis accelerometer data and tri-axis gyroscopic data (Zhang and Sawchuk, 2012). RWC comprises underwater acoustic recordings used to detect North Atlantic right whale calls in the presence of ocean noise Vickers and Middlehurst . TEE originates from the FORTE satellite and classifies transient electromagnetic events (e.g., lightning strokes) based on waveform signatures Eads et al. (2002).
@@ Appendix B Addressing Probing Concerns
Next, we follow the guidance of Belinkov et al. (2017a) and Conneau et al. (2018), we train a classifier on an LLM with randomly initialized weights. Specifically, we chose Llama-3.2-11B-Vision-Instruct as it can be directly compared to its fine-tuned version.
@@
As an upper-bound reference, established baselines designed specifically to learn discriminative time-series representations, following prior probing work that contrasts frozen representations with task-optimized or state-of-the-art models Belinkov et al. (2017b); Conneau et al. (2018); Liu et al. (2019).
@@ Appendix E Hyperparameters
For reproducibility, we provide the complete hyperparameter configurations for all baseline methods. These settings were either adopted from the original publications or carefully tuned to ensure fair comparison across all methods on our datasets.
@@ E.2 Attend
We implement our experiments using PyTorch. Our network is trained end-to-end using the Adam optimizer Adam and others (2014) with a learning rate that decays periodically. The training data is partitioned into segments using a sliding window approach with 50% overlap between adjacent windows. We employ mixup augmentation as a regularization technique and incorporate center loss to encourage intra-class compactness in the learned feature space. Dropout is applied at multiple stages of the network, including the RNN layers and the classifier, to prevent overfitting. Model performance is evaluated using 5-fold cross-validation to ensure robust and reliable results. All hyperparameters are summarized in Table 8.

### v2

@@ 1 Introduction
LLMs demonstrate broad generalization across natural language tasks, supporting flexible pattern matching, contextual reasoning, and symbolic manipulation. These capabilities have motivated efforts to extend LLMs to non-linguistic data Fons et al. (2024); Mai et al. (2024). Recent work has explored LLMs for time-series forecasting, classification, and analysis Jin et al. (2024a); Gruver et al. (2023); Tan et al. (2024); Morales-García et al. (2025), often relying on major modifications and retraining. However, strong end-to-end performance in these settings does not necessarily imply that LLMs encode discriminative temporal structure, as results may instead reflect task formulation, additional fine-tuning, or prompting choices.
@@
Merrill et al. (2024) show that, despite some success in zero-shot forecasting, LLMs struggle to reliably extract discriminative information from time series, performing near-randomly on reasoning and question-answering tasks that require interpreting temporal structure. Others similarly show that removing the LLM component from time-series pipelines often does not degrade performance and can even improve it, calling into question the necessity of language models for these tasks Tan et al. (2024). In contrast, a growing body of work argues that LLM-centric approaches may enable more general and interactive forms of time series understanding that extend beyond task-specific prediction Jin et al. (2024b). Together, these mixed findings raise a central question: do observed limitations reflect a lack of time-series understanding in LLMs, or a mismatch between what these models represent and how that information is extracted?
@@ 2 Related Work
Probing and Hidden Representations in LLMs. Early work on Layer-wise Relevance Propagation (LRP) introduced methods for attributing model predictions to intermediate representations Bach et al. (2015) which later extended to more complex architectures Binder et al. (2016); Arras et al. (2017). Building on this work, probing methods emerged to explicitly measure what information these representations encode by testing their linear separability with lightweight classifiers Alain and Bengio (2018). In recent years, probing methods have increasingly been used as to extract features for a wide range of downstream applications. Zhu et al. (2025) and Wang et al. (2025), show that hidden states can be leveraged to estimate the difficulty an LLM is anticipated to have on a query prior to generation. More broadly, prior work has demonstrated that hidden representations encode rich and actionable signals that can be exploited across diverse settings, including alignment Kong et al. (2024); Li et al. (2023); Zhang et al. (2025), safety Wang et al. (2024); Lu et al. (2025), interpretability Ghandeharioun et al. (2024); Jacobi and Niv (2025), and robustness Lad et al. (2024); Yan et al. (2024).
@@
Time-Series Research. Time series research spans a wide range of objectives, with forecasting and imputation forming the historical foundation of the field. Classical statistical methods, such as ARIMA and Kalman filters, model linear temporal dependencies Box and Jenkins (1990); Hamilton (1994); Hyndman and Athanasopoulos (2018), while neural architectures, including LSTMs, convolutional networks, and transformers, enable richer nonlinear and long-range temporal modeling Hochreiter and Schmidhuber (1997); Bai et al. (2018); Zhou et al. (2021); Wu et al. (2021); Lim et al. (2021).
Building on this progression, recent work explores how large language models can be adapted for time series analysis. Existing approaches largely follow a small number of recurring patterns Zhang et al. (2024): direct prompting that treats time series as text, often with modified numerical tokenization Gruver et al. (2023); Xue and Salim (2024); Liu et al. (2023); Xie et al. (2023); discretization or quantization schemes that map signals to token sequences Chung et al. (2023); Duan et al. (2023); Yang et al. (2024); Rubenstein et al. (2023); Yu et al. (2023); and architectures that introduce dedicated time-series encoders aligned with language model representations King et al. (2023); Sun et al. (2024); Zhou et al. (5555); Qiu et al. (2023); Zhou et al. (2023); Chang et al. (2025); Jin et al. (2024a). Related work also explores visual time-series representations processed by vision–language models or uses LLMs as controllers for external analysis tools Girdhar et al. (2023); Su et al. (2023); Moon et al. (2023); Zhang et al. (2023).
@@ A.1 Baselines
Attend follows the dedicated time-series encoder family. It uses self-attention to model cross-channel and temporal dependencies in multivariate sensor data and learns discriminative representations directly from raw time series, without adapting language-model token spaces Abedin et al. (2021). InstructTime belongs to the time-series quantization and instruction-based prompting family. It discretizes time series into vector-quantized tokens, aligns them with domain-specific textual instructions, and uses generative instruction tuning to produce label text via next-token prediction Cheng et al. (2025). OneFitsAll follows the encoder-alignment family. It reuses a frozen pretrained transformer from language or vision domains and learns lightweight input projections and normalization layers to map time-series patches into the model’s embedding space, enabling a single backbone across tasks Zhou et al. (2023). TS2Vec represents dedicated time-series representation learning. It learns multi-scale contextual embeddings through hierarchical contrastive learning over augmented temporal views, which are aggregated and used by simple downstream classifiers without language models Yue et al. (2022). Moment belongs to the TS-Encoder family. It is a high-capacity transformer encoder pre-trained on large-scale, multi-domain time series using masked time-series modeling, and serves as a general-purpose representation learner that can be adapted to classification, forecasting, anomaly detection, and imputation via lightweight task-specific heads or linear probing (Goswami et al., 2024).
@@ A.2 Models
To assess the generality of our findings, we evaluate a diverse set of pretrained models spanning different architectures and scales. [^1] All models are open source, as access to internal hidden states is required for probing. Motivated by prior work showing that visualizations provide a strong representation for time series classification Liu et al. (2025), we primarily focus on vLLMs. Specifically, we evaluate Llama-3.2-11B-Vision-Instruct Grattafiori et al. (2024), Qwen/Qwen2.5-VL-32B-Instruct Bai et al. (2025), and Mistral-Small-3.1-24B-Instruct-2503 MistralAI . To verify that our conclusions extend beyond multimodal architectures, we also evaluate two text-only LLMs: GPT-oss OpenAI et al. (2025) and google/gemma-2-9b-it Team et al. (2024).
@@ A.3 Datasets
CTU records 24-hour household electricity usage patterns and classifies computers as desktops or laptops, capturing differences in device power-consumption cycles Chen et al. (2015). EMG consists of muscle response recordings under electrical stimulation, categorized into healthy, myopathy, and neuropathy classes Goldberger et al. (2000). HAR contains multivariate accelerometer recordings of human activities such as walking, sitting, and stair climbing, and exhibits substantial inter-subject variability Anguita et al. (2013). HAD similarly contains multivariate human activity data; we use both the tri-axis accelerometer data and tri-axis gyroscopic data (Zhang and Sawchuk, 2012). RWC comprises underwater acoustic recordings used to detect North Atlantic right whale calls in the presence of ocean noise Cox et al. (2006). TEE originates from the FORTE satellite and classifies transient electromagnetic events (e.g., lightning strokes) based on waveform signatures Eads et al. (2002).
@@ Appendix B Addressing Probing Concerns
Next, we follow the guidance of Belinkov et al. (2017) and Conneau et al. (2018), we train a classifier on an LLM with randomly initialized weights. Specifically, we chose Llama-3.2-11B-Vision-Instruct as it can be directly compared to its fine-tuned version.
@@
As an upper-bound reference, established baselines designed specifically to learn discriminative time-series representations, following prior probing work that contrasts frozen representations with task-optimized or state-of-the-art models Belinkov et al. (2017); Conneau et al. (2018); Liu et al. (2019).
@@ Appendix E Hyperparameters
wFor reproducibility, we provide the complete hyperparameter configurations for all baseline methods. These settings were either adopted from the original publications or carefully tuned to ensure fair comparison across all methods on our datasets.
@@ E.2 Attend
We implement our experiments using PyTorch. Our network is trained end-to-end using the Adam optimizer Kingma and Ba (2017) with a learning rate that decays periodically. The training data is partitioned into segments using a sliding window approach with 50% overlap between adjacent windows. We employ mixup augmentation as a regularization technique and incorporate center loss to encourage intra-class compactness in the learned feature space. Dropout is applied at multiple stages of the network, including the RNN layers and the classifier, to prevent overfitting. Model performance is evaluated using 5-fold cross-validation to ensure robust and reliable results. All hyperparameters are summarized in Table 8.